If a solid PQ have NaCl type structure is heated so that the ions along one body diagonal are lost and trivalent ions(R) enter interstitial sites to maintain the electrical neutrality so that the new empirical formula unit is PxQyRz(x,y,z are smallest possible integers). The values of x+y+z is [Q- form the lattice]

1
Expert's answer
2021-06-07T03:14:34-0400

"x+y+c=4+2\u22121=5"